# Minimax estimation of functionals in sparse vector model with correlated observations

### v1

@@ 5 Optimal testing for signal detection
${\mathcal{R}}^{*}(s,d,\rho,\bar{r}_{\gamma})\geq\exp(1-\exp(4c_{4}\gamma^{2}))/2.$
For any $\eta\in(0,1/2)$, the equation $\exp(1-\exp(4c_{4}\gamma^{2}))/2=\eta$ has the solution
$\gamma_{\eta}=\sqrt{(4c_{4})^{-1}\log\log\big{(}\frac{e}{2\eta}\big{)}}.$
@@
The results of this section cannot be directly compared to the prior work on sparse testing discussed in the Introduction since those papers were analyzing minimax optimality for a fixed matrix ${\boldsymbol{\Sigma}}$. In their settings, $H_{0}$ was a simple hypothesis (with a given matrix ${\boldsymbol{\Sigma}}$, which was assumed diagonal, except for [15, 12]) while here we deal with minimax optimality over classes of matrices ${\boldsymbol{\Sigma}}$, and in our case $H_{0}$ is a composite hypothesis. Of course, for specific fixed matrices in the class one gets, in general, better rates than those achieved uniformly over a class. This can be seen in the setting with equicorrelation matrix detailed in [12]. In [15, Theorem 7] the authors provide the minimax rate for a testing problem in a change-point framework, which implies the minimax rate of the order $\sqrt{\|{\boldsymbol{\Sigma}}\|_{F}}$ for testing in model (1) under the no sparsity scenario $s=d$ with fixed positive definite matrix ${\boldsymbol{\Sigma}}$ and $\sigma=1$. Note that it agrees with our conclusions for this particular setting. Indeed, when ${\boldsymbol{\Sigma}}$ is fixed with entries independent of $d$ we have $\|{\boldsymbol{\Sigma}}\|_{F}\lesssim d$, so that for $s=d$ our Theorem 1 gives the rate $\psi(s,{\boldsymbol{\Sigma}})=\|{\boldsymbol{\Sigma}}\|_{F}$ for the squared loss leading to the $\sqrt{\|{\boldsymbol{\Sigma}}\|_{F}}$ rate of testing for this particular case.
@@ 6.1 Properties of correlated Gaussian variables
There exists an absolute constant $C>0$ such that the following statements hold for two standard Gaussian random variables $\zeta$ and $\eta$ with correlation coefficient $\nu$.
@@
    $\mathbf{Cov}[(\zeta^{2}-\beta)\mathbbm{1}(|\zeta|>\tau),(\eta^{2}-\beta)\mathbbm{1}(|\eta|>\tau)]\leq C\nu^{2}\tau^{4}\exp(-\tau^{2}/2).$
@@
    $|\mathbf{Cov}[\mathbbm{1}(\zeta^{2}\leq\tau),\mathbbm{1}(\eta^{2}\leq\tau)]|\leq C\nu^{2}(\tau^{2}+1)e^{-\tau/3}.$
@@
    $|\mathbf{Cov}[\cos(t(\mu_{1}+\sigma_{1}\zeta)),\cos(t(\mu_{2}+\sigma_{2}\eta))]|\leq C\big{(}|\nu^{2}\cos(t\mu_{1})\cos(t\mu_{2})|+|\nu\sin(t\mu_{1})\sin(t\mu_{2})|\big{)}.$
@@
    $|\mathbf{Cov}[\mathbbm{1}(\zeta^{2}\leq\tau),\mathbbm{1}(\eta^{2}\leq\tau)]|\leq\mathbf{Var}[\mathbbm{1}(\eta^{2}\leq\tau)]\leq\mathbf{P}[\eta^{2}>\tau]\leq C\nu^{2}e^{-\tau/2},$
@@
[…]
    where $I_{1}$ is the interval with endpoints $\frac{\sqrt{\tau}-\nu u}{\sqrt{1-\nu^{2}}}$ and $\frac{\sqrt{\tau}+\nu u}{\sqrt{1-\nu^{2}}}$, $I_{2}=[\sqrt{\tau},\frac{\sqrt{\tau}}{\sqrt{1-\nu^{2}}}]$, and we have used the triangle inequality and the relation between divided differences and derivatives (see, e.g., [16, (2.1.4.3)]). Since $|\nu|\leq\frac{1}{5}$ and $|u|\leq\sqrt{\tau}$, we have $I_{1}\subseteq[\sqrt{\frac{2\tau}{3}},\sqrt{\frac{3\tau}{2}}]$. Note that $\Phi^{\prime}(x)=\frac{1}{\sqrt{2\pi}}e^{-x^{2}/2}$ and $\Phi^{\prime\prime}(x)=-\frac{x}{\sqrt{2\pi}}e^{-x^{2}/2}$, and $\frac{1}{\sqrt{1-\nu^{2}}}-1=\frac{\nu^{2}}{\sqrt{1-\nu^{2}}+1-\nu^{2}}$. Therefore, the upper bound (32) yields
    $\left|\Phi\left(\frac{\sqrt{\tau}-\nu u}{\sqrt{1-\nu^{2}}}\right)-\Phi\left(\frac{-\sqrt{\tau}-\nu u}{\sqrt{1-\nu^{2}}}\right)-\Phi(\sqrt{\tau})+\Phi(-\sqrt{\tau})\right|\leq C^{\prime}\nu^{2}(\tau^{3/2}+\tau^{1/2})e^{-\tau/3},$
@@
    $=\frac{1}{2}{\mathbf{E}}[\cos t(\tilde{\zeta}+\tilde{\eta})]+\frac{1}{2}{\mathbf{E}}[\cos t(\tilde{\zeta}-\tilde{\eta})]-{\mathbf{E}}[\cos(t\tilde{\zeta})]{\mathbf{E}}[\cos(t\tilde{\eta})]$ (35)
@@
    For $|\nu|\geq\frac{1}{2}$ these upper bounds hold trivially. For $|\nu|\leq\frac{1}{2}$, using the fact that $|\delta|\leq|\nu|\leq\frac{1}{2}$ and applying the relation between divided differences and derivatives (see, e.g., [16, (2.1.4.3)] we obtain
@@
    $\mathbf{Var}\left[\sum_{i\in\mathcal{U}}\mathbbm{1}(\varepsilon_{i}^{2}\leq\sigma_{i}^{2}\tau)\right]\leq C\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}^{2}(\tau^{2}+1)e^{-\tau/3},$
@@
    $\mathbf{Var}\left[\sum_{i=1}^{d}\cos(t(\mu_{i}+\varepsilon_{i}))\right]\leq C\left(\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}^{2}+s\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}\right).$
@@
$\mathbf{Var}\left[\sum_{i\in\mathcal{U}}\varepsilon_{i}^{2}\mathbbm{1}(|\varepsilon_{i}|>\sigma_{i}\tau)\right]\leq 2\mathbf{Var}\left[\sum_{i\in\mathcal{U}}(\varepsilon_{i}^{2}-\sigma_{i}^{2}\beta)\mathbbm{1}(|\varepsilon_{i}|>\sigma_{i}\tau)\right]+2\mathbf{Var}\left[\sum_{i\in\mathcal{U}}\sigma_{i}^{2}\beta\mathbbm{1}(|\varepsilon_{i}|>\sigma_{i}\tau)\right].$
@@ 6.2.1 Proof of Theorem 1
$\hat{Q}-\|{\boldsymbol{\theta}}\|_{2}^{2}=\sum_{i=1}^{d}Y_{i}^{2}-\sigma^{2}\sigma_{i}^{2}-\theta_{i}^{2}=\sigma^{2}\sum_{i=1}^{d}(\varepsilon_{i}^{2}-\sigma_{i}^{2})+2\sigma\sum_{i\in\mathcal{S}}\theta_{i}\varepsilon_{i},$
@@
$\qquad+\sum_{i\in\mathcal{S}}\sigma^{2}(\varepsilon_{i}^{2}-\sigma_{i}^{2}\beta_{s})\mathbbm{1}(|\varepsilon_{i}|\leq\sigma\tau)-\sum_{i\in\mathcal{S}}(Y_{i}^{2}-\sigma^{2}\sigma_{i}^{2}\beta_{s})\mathbbm{1}(|Y_{i}|\leq\sigma\sigma_{i}\tau).$ (54)
@@
[…]
@@
$\left|\sum_{i\in\mathcal{S}}(\varepsilon_{i}^{2}-\sigma^{2}\sigma_{i}^{2}\beta_{s})\mathbbm{1}(|\varepsilon_{i}|\leq\sigma\sigma_{i}\tau)-\sum_{i\in\mathcal{S}}(Y_{i}^{2}-\sigma^{2}\sigma_{i}^{2}\beta_{s})\mathbbm{1}(|Y_{i}|\leq\sigma\sigma_{i}\tau)\right|\lesssim s\sigma^{2}\tau^{2}\max_{i\in\mathcal{S}}\sigma_{i}^{2}.$
@@
${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[(\hat{Q}-\|{\boldsymbol{\theta}}\|_{2}^{2})^{2}\right]\lesssim\sigma^{2}\|{\boldsymbol{\theta}}\|_{2}^{2}\lambda_{\max}({\boldsymbol{\Sigma}}_{\mathcal{S},\mathcal{S}})+\sigma^{4}\tau^{4}\left(s^{2}+\|{\boldsymbol{\Sigma}}\|_{F}^{2}e^{-\tau^{2}/2}\right).$
@@
$\lambda_{\max}({\boldsymbol{\Sigma}}_{\mathcal{S},\mathcal{S}})\leq\lambda_{\max}({\boldsymbol{\Sigma}})\wedge\|{\boldsymbol{\Sigma}}_{\mathcal{S},\mathcal{S}}\|_{F}\leq\lambda_{\max}({\boldsymbol{\Sigma}})\wedge s\max_{i\in\mathcal{S}}\sigma_{i}^{2}\leq\lambda_{\max}({\boldsymbol{\Sigma}})\wedge s.$
@@ 6.2.2 Proof of Theorem 2
Let $\gamma>0$ be such that $w(\gamma)>0$. Since $w(\cdot)$ is a monotone non-decreasing function, for any estimator $\hat{T}$ we have
$\sup_{({\boldsymbol{\theta}},{\boldsymbol{\Sigma}})\in\Theta(s,d)}{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\,w\big{(}\sigma^{-1}\psi(s,{\boldsymbol{\Sigma}})^{-1/2}|\hat{T}-\|{\boldsymbol{\theta}}\|_{2}|\big{)}\geq w(\gamma)\sup_{({\boldsymbol{\theta}},{\boldsymbol{\Sigma}})\in\Theta(s,d)}{\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\big{(}|\hat{T}-\|{\boldsymbol{\theta}}\|_{2}|\geq\gamma\sigma\sqrt{\psi(s,{\boldsymbol{\Sigma}})}\big{)}$ (64)
$\qquad\qquad\geq\frac{w(\gamma)}{2}\left[\mathbf{P}_{\mathbf{0},{\boldsymbol{I}}}\big{(}|\hat{T}|\geq\gamma\sigma\sqrt{\psi(s,{\boldsymbol{I}})}\big{)}+\sup_{{\boldsymbol{\theta}}\in B_{0}(s,d):\|{\boldsymbol{\theta}}\|_{2}\geq 2\gamma\sigma\sqrt{\psi(s,{\boldsymbol{I}})}}\mathbf{P}_{{\boldsymbol{\theta}},{\boldsymbol{I}}}\big{(}|\hat{T}|\leq\gamma\sigma\sqrt{\psi(s,{\boldsymbol{I}})}\big{)}\right]$ (65)
$\qquad\qquad\geq\frac{w(\gamma)}{2}\inf_{\Psi}\left[\mathbf{P}_{\mathbf{0},{\boldsymbol{I}}}(\Psi=1)+\sup_{{\boldsymbol{\theta}}\in B_{0}(s,d):\|{\boldsymbol{\theta}}\|_{2}\geq 2\gamma\sigma\sqrt{\phi(s,d)}}\mathbf{P}_{{\boldsymbol{\theta}},{\boldsymbol{I}}}(\Psi=0)\right]$ (66)
where $\inf_{\Psi}$ denotes the infimum over all tests and we have used the fact that $\psi(s,{\boldsymbol{I}})=\phi(s,d)$. To complete the proof, it remains to show that
$\inf_{\Psi}\left[\mathbf{P}_{\mathbf{0},{\boldsymbol{I}}}(\Psi=1)+\sup_{{\boldsymbol{\theta}}\in B_{0}(s,d):\|{\boldsymbol{\theta}}\|_{2}\geq 2\gamma\sigma\sqrt{\phi(s,d)}}\mathbf{P}_{{\boldsymbol{\theta}},{\boldsymbol{I}}}(\Psi=0)\right]\geq c(\gamma)>0$ (67)
for all $\gamma>0$, where $c(\gamma)$ depends only on $\gamma$. This is done following the same lines as in the proof of [7, Theorem 12]. Note that [7, Theorem 12] obtains (67) for sufficiently small $\gamma$, while we need it here for all $\gamma>0$. Extending the proof in [7] to all $\gamma>0$ is straightforward by using the inequality $V(P,Q)\leq 1-\exp(-\chi^{2}(P,Q))/2$ instead of $V(P,Q)\leq\sqrt{\chi^{2}(P,Q)}$, where $V(P,Q)$ and $\chi^{2}(P,Q)$ are the total variation distance and the chi-square divergence between two probability measures $P$ and $Q$ (cf. [17, Theorem 2.2(iii)]). Following the lines of [7, Theorem 12] with this modification we obtain that the bound (67) holds with $c(\gamma)=\exp(1-\exp(4\gamma^{2}))/2$.
6.2.3 Proof of Theorem 3
@@
$R^{*}(s,d,\rho,\phi):=\inf_{\hat{T}}\sup_{({\boldsymbol{\theta}},{\boldsymbol{\Sigma}})\in\Theta(s,d,\rho)}{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}[w(\sigma^{-1}\phi^{-1/2}|\hat{T}-\|{\boldsymbol{\theta}}\|_{2}|)].$
@@
$\tilde{R}^{*}(s,d,\phi):=\inf_{\hat{T}}\sup_{{\boldsymbol{\theta}}\in B_{0}(s,d),{\boldsymbol{\Sigma}}={\boldsymbol{I}}_{d\times d}}{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}[w(\sigma^{-1}\phi^{-1/2}|\hat{T}-\|{\boldsymbol{\theta}}\|_{2}|)].$
@@
$\tilde{\Theta}:=\left\{(\underset{r\;\mathrm{entries}}{\underbrace{\tilde{\theta}_{1},\cdots,\tilde{\theta}_{1}}},\cdots,\underset{r\;\mathrm{entries}}{\underbrace{\tilde{\theta}_{p},\cdots,\tilde{\theta}_{p}}},0,\cdots,0)^{\top}\in\mathbb{R}^{d}:\sum_{i=1}^{p}\mathbbm{1}(\tilde{\theta}_{i}\neq 0)\leq 1\right\}.$

### v2

@@ 5 Optimal testing for signal detection
${\mathcal{R}}^{*}(s,d,\rho,\bar{r}_{\gamma})\geq\exp(1-\exp(c_{4}\gamma^{2}))/2.$ (46)
For any $\eta\in(0,1/2)$, the equation $\exp(1-\exp(c_{4}\gamma^{2}))/2=\eta$ has the solution
$\gamma_{\eta}=\sqrt{(c_{4})^{-1}\log\log\big(\frac{e}{2\eta}\big)}.$
@@
The results of this section deal with minimax optimality on a class of matrices, while the prior work on sparse testing discussed in the Introduction was analyzing minimax optimality for a fixed matrix $\boldsymbol{\Sigma}$}. In those papers, $H_{0}$ was a simple hypothesis (with a given matrix $\boldsymbol{\Sigma}$}, which was assumed diagonal, except for [10, 13, 16]), while here we deal with minimax optimality over classes of matrices $\boldsymbol{\Sigma}$}, and in our case $H_{0}$ is a composite hypothesis. Of course, for specific fixed matrices in the class one gets, in general, better rates than those achieved uniformly over a class. This can be seen in the setting with equicorrelation matrix detailed in Kotekal and Gao [13], and in Hall and Jin [10] for a different instance of sparse testing problem in asymptotics as $d\to\infty$.
In [16, Theorem 7] the authors provide the minimax rate for a testing problem in a change-point framework, which implies the minimax rate of the order $\sqrt{\|{\boldsymbol{\Sigma}}\|_{F}}$ for testing in model (2) under the no sparsity scenario $s=d$ with fixed positive definite matrix $\boldsymbol{\Sigma}$} and $\sigma=1$. Note that it agrees with our conclusions for this particular setting. Indeed, when $\boldsymbol{\Sigma}$} is fixed with entries independent of $d$ we have $\|{\boldsymbol{\Sigma}}\|_{F}\lesssim d$, so that for $s=d$ our Theorem 1 gives the rate $\psi(s,{\boldsymbol{\Sigma}})=\|{\boldsymbol{\Sigma}}\|_{F}$ for the squared loss leading to the $\sqrt{\|{\boldsymbol{\Sigma}}\|_{F}}$ rate of testing for this particular case.
Remark 1.
An interesting phenomenon described in Hall and Jin [10] and further noticed in other context by Kotekal and Gao [13] consists in the fact that, for some specific families of matrices $\boldsymbol{\Sigma}$}, the minimax rate of testing can be better than in the case of independent noise, and the hardest subproblem corresponds to the identity covariance matrix. In contrast, in our setting this is not the case. Indeed, inspection of the proof shows that our lower bound is obtained on correlated data that correspond to block-diagonal matrices and not to the identity covariance matrix. The independent case appears in the proof only as a technical element – after a suitable reduction we apply the lower bounds for the independent case from Collier et al. [7]. Our contribution is a reduction argument that leads to a contraction of the original problem to testing two composite hypotheses on a smaller scale. Importantly, the reduction step itself does not rely on the hypotheses being composite. The intuition is that, with a stronger correlation on $\boldsymbol{\Sigma}$}, the effective sample size of the noise vector $\boldsymbol{\varepsilon}$} becomes smaller, with the effective sample size being directly governed by the parameter $\rho$. In our lower bound construction, this reduction is realized through a block-wise covariance structure, where the block size is also determined by $\rho$. Consequently, the estimation error of the mean parameter within each block is amplified by the block size. This construction provides a convenient and analytically tractable instance that captures the essential difficulty of the problem.
@@ 6.1 Properties of correlated Gaussian variables
There exists an absolute constant $C>0$ such that the following statements hold for two jointly normal standard Gaussian random variables $\zeta$ and $\eta$ that are jointly normally distributed with correlation coefficient $\nu$.
@@
    $\mathbf{Cov}[(\zeta^{2}-\beta)\mathbbm{1}(|\zeta|>\tau),(\eta^{2}-\beta)\mathbbm{1}(|\eta|>\tau)]\leq C\nu^{2}\tau^{4}\exp(-\tau^{2}/2).$ (48)
@@
    $|\mathbf{Cov}[\mathbbm{1}(\zeta^{2}\leq\tau),\mathbbm{1}(\eta^{2}\leq\tau)]|\leq C\nu^{2}(\tau^{2}+1)e^{-\tau/3}.$ (49)
@@
    $|\mathbf{Cov}[\cos(t(\mu_{1}+\sigma_{1}\zeta)),\cos(t(\mu_{2}+\sigma_{2}\eta))]|\leq C\big(|\nu^{2}\cos(t\mu_{1})\cos(t\mu_{2})|+|\nu\sin(t\mu_{1})\sin(t\mu_{2})|\big).$ (50)
@@
    $|\mathbf{Cov}[\mathbbm{1}(\zeta^{2}\leq\tau),\mathbbm{1}(\eta^{2}\leq\tau)]|\leq\mathbf{Var}[\mathbbm{1}(\eta^{2}\leq\tau)]\leq\mathbf{P}[\eta^{2}>\tau]\leq C\nu^{2}e^{-\tau/2},$ (55)
@@
    $=\penalty 10000\$$\left|\Phi\left(\frac{\sqrt{\tau}+\nu u}{\sqrt{1-\nu^{2}}}\right)+\Phi\left(\frac{\sqrt{\tau}-\nu u}{\sqrt{1-\nu^{2}}}\right)-2\Phi\left(\frac{\sqrt{\tau}}{\sqrt{1-\nu^{2}}}\right)+2\Phi\left(\frac{\sqrt{\tau}}{\sqrt{1-\nu^{2}}}\right)-2\Phi\left(\sqrt{\tau}\right)\right|$ (60)
@@
    where $I_{1}$ is the interval with endpoints $\frac{\sqrt{\tau}-\nu u}{\sqrt{1-\nu^{2}}}$ and $\frac{\sqrt{\tau}+\nu u}{\sqrt{1-\nu^{2}}}$, $I_{2}=[\sqrt{\tau},\frac{\sqrt{\tau}}{\sqrt{1-\nu^{2}}}]$, and we have used the triangle inequality and the relation between divided differences and derivatives (see, e.g., [17, (2.1.4.3)]). Since $|\nu|\leq\frac{1}{5}$ and $|u|\leq\sqrt{\tau}$, we have $I_{1}\subseteq[\sqrt{\frac{2\tau}{3}},\sqrt{\frac{3\tau}{2}}]$. Note that $\Phi^{\prime}(x)=\frac{1}{\sqrt{2\pi}}e^{-x^{2}/2}$ and $\Phi^{\prime\prime}(x)=-\frac{x}{\sqrt{2\pi}}e^{-x^{2}/2}$, and $\frac{1}{\sqrt{1-\nu^{2}}}-1=\frac{\nu^{2}}{\sqrt{1-\nu^{2}}+1-\nu^{2}}$. Therefore, the upper bound (61) yields
    $\left|\Phi\left(\frac{\sqrt{\tau}-\nu u}{\sqrt{1-\nu^{2}}}\right)-\Phi\left(\frac{-\sqrt{\tau}-\nu u}{\sqrt{1-\nu^{2}}}\right)-\Phi(\sqrt{\tau})+\Phi(-\sqrt{\tau})\right|\leq C^{\prime}\nu^{2}(\tau^{3/2}+\tau^{1/2})e^{-\tau/3},$ (62)
@@
[…]
@@
    For $|\nu|\geq\frac{1}{2}$ these upper bounds hold trivially. For $|\nu|\leq\frac{1}{2}$, using the fact that $|\delta|\leq|\nu|\leq\frac{1}{2}$ and applying the relation between divided differences and derivatives (see, e.g., [17, (2.1.4.3)] we obtain
@@
    $\mathbf{Var}\left[\sum_{i\in\mathcal{U}}\mathbbm{1}(\varepsilon_{i}^{2}\leq\sigma_{i}^{2}\tau)\right]\leq C\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}^{2}(\tau^{2}+1)e^{-\tau/3},$ (72)
@@
    $\mathbf{Var}\left[\sum_{i=1}^{d}\cos(t(\mu_{i}+\varepsilon_{i}))\right]\leq C\left(\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}^{2}+s\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}\right).$ (73)
@@
$\mathbf{Var}\left[\sum_{i\in\mathcal{U}}\varepsilon_{i}^{2}\mathbbm{1}(|\varepsilon_{i}|>\sigma_{i}\tau)\right]\leq 2\mathbf{Var}\left[\sum_{i\in\mathcal{U}}(\varepsilon_{i}^{2}-\sigma_{i}^{2}\beta)\mathbbm{1}(|\varepsilon_{i}|>\sigma_{i}\tau)\right]+2\mathbf{Var}\left[\sum_{i\in\mathcal{U}}\sigma_{i}^{2}\beta\mathbbm{1}(|\varepsilon_{i}|>\sigma_{i}\tau)\right].$ (76)
@@ 6.2.1 Proof of Theorem 1
$\hat{Q}-\|{\boldsymbol{\theta}}\|_{2}^{2}=\sum_{i=1}^{d}Y_{i}^{2}-\sigma^{2}\sigma_{i}^{2}-\theta_{i}^{2}=\sigma^{2}\sum_{i=1}^{d}(\varepsilon_{i}^{2}-\sigma_{i}^{2})+2\sigma\sum_{i\in\mathcal{S}}\theta_{i}\varepsilon_{i},$ (82)
@@
$\qquad+\sum_{i\in\mathcal{S}}\sigma^{2}(\varepsilon_{i}^{2}-\sigma_{i}^{2}\beta_{s})\mathbbm{1}(|\varepsilon_{i}|\leq\sigma_{i}\tau)-\sum_{i\in\mathcal{S}}(Y_{i}^{2}-\sigma^{2}\sigma_{i}^{2}\beta_{s})\mathbbm{1}(|Y_{i}|\leq\sigma\sigma_{i}\tau).$ (88)
@@
$\beta_{s}=1+\frac{\tau\frac{1}{\sqrt{2\pi}}e^{-\tau^{2}/2}}{\int_{\tau}^{\infty}\frac{1}{\sqrt{2\pi}}e^{-x^{2}/2}dx}<1+\frac{\tau(\tau+\sqrt{\tau^{2}+4})}{2}\leq\tau^{2}+\tau+1,$ (89)
@@
$\left|\sum_{i\in\mathcal{S}}(\varepsilon_{i}^{2}-\sigma^{2}\sigma_{i}^{2}\beta_{s})\mathbbm{1}(|\varepsilon_{i}|\leq\sigma_{i}\tau)-\sum_{i\in\mathcal{S}}(Y_{i}^{2}-\sigma^{2}\sigma_{i}^{2}\beta_{s})\mathbbm{1}(|Y_{i}|\leq\sigma\sigma_{i}\tau)\right|\lesssim s\sigma^{2}\tau^{2}\max_{i\in\mathcal{S}}\sigma_{i}^{2}.$ (90)
@@
${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[(\hat{Q}-\|{\boldsymbol{\theta}}\|_{2}^{2})^{2}\right]\lesssim\sigma^{2}\|{\boldsymbol{\theta}}\|_{2}^{2}\lambda_{\max}({\boldsymbol{\Sigma}}_{\mathcal{S},\mathcal{S}})+\sigma^{4}\tau^{4}\left(s^{2}+\|{\boldsymbol{\Sigma}}\|_{F}^{2}e^{-\tau^{2}/2}\right).$ (93)
@@
$\lambda_{\max}({\boldsymbol{\Sigma}}_{\mathcal{S},\mathcal{S}})\leq\lambda_{\max}({\boldsymbol{\Sigma}})\wedge\|{\boldsymbol{\Sigma}}_{\mathcal{S},\mathcal{S}}\|_{F}\leq\lambda_{\max}({\boldsymbol{\Sigma}})\wedge s\max_{i\in\mathcal{S}}\sigma_{i}^{2}\leq\lambda_{\max}({\boldsymbol{\Sigma}})\wedge s.$ (94)
@@ 6.2.2 Proof of Theorem 2
$R^{*}(s,d,\rho,\phi):=\inf_{\hat{T}}\sup_{({\boldsymbol{\theta}},{\boldsymbol{\Sigma}})\in\Theta(s,d,\rho)}{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}[w(\sigma^{-1}\phi^{-1/2}|\hat{T}-\|{\boldsymbol{\theta}}\|_{2}|)].$ (101)
@@
$\tilde{R}^{*}(s,d,\phi):=\inf_{\hat{T}}\sup_{{\boldsymbol{\theta}}\in B_{0}(s,d),{\boldsymbol{\Sigma}}={\boldsymbol{I}}_{d\times d}}{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}[w(\sigma^{-1}\phi^{-1/2}|\hat{T}-\|{\boldsymbol{\theta}}\|_{2}|)].$ (102)
@@
$\tilde{\Theta}:=\left\{(\underset{r\;\mathrm{entries}}{\underbrace{\tilde{\theta}_{1},\cdots,\tilde{\theta}_{1}}},\cdots,\underset{r\;\mathrm{entries}}{\underbrace{\tilde{\theta}_{p},\cdots,\tilde{\theta}_{p}}},0,\cdots,0)^{\top}\in\mathbb{R}^{d}:\sum_{i=1}^{p}\mathbbm{1}(\tilde{\theta}_{i}\neq 0)\leq 1\right\}.$ (107)